OpenStudy (anonymous):

Rationalize the denominators of these expressions to place in the form sqrt(a)/b: 1/sqrt(7) sqrt(2/7)

OpenStudy (anonymous):

To rationalize the denominators, multiply the top and bottom by the sqrt on the bottom.

OpenStudy (anonymous):

\[\frac{1}{\sqrt{7}}=\frac{1}{\sqrt{7}}\times \frac{\sqrt{7}}{\sqrt{7}}=\frac{\sqrt{7}}{7}\]

OpenStudy (anonymous):

sqrt{2}/2

OpenStudy (anonymous):

\[\sqrt{\frac{2}{7}}=\frac{\sqrt{2}}{\sqrt{7}}=\frac{\sqrt{2}}{\sqrt{7}}\times \frac{\sqrt{7}}{\sqrt{7}}=\frac{\sqrt{14}}{7}\]\]

OpenStudy (anonymous):

time for lunch